A ball of mass 5g is thrown downwards with velocity 10mps from height of 19.5m it penetrates soil by 50cm , the change in mechanical energy is

Grade:12

1 Answers

Ayushi
3 years ago
U= 10mps  m= 5g = 5/1000 kg  h= 19.5 m 
Kinetic energy = 1/2 mv²
= 1/2×5/1000×10×10 
=1/4 J 
Potential energy = mg(h1+h2) 
= 5/1000 ×10 ( 19.5+50/100)
= 1J 
Hence, M.E = KE +U 
                    = 1/4 + 1 = 1.25 J